OpenStudy (anonymous):

write an equation of the line containing the given points and perpendicular to the given line. use form y=mx+b (5,9);5x+y=8

OpenStudy (anonymous):

THIS LINE HAS A GRADIENT of 5 The perpendicular would therefore have a gradient of -1/5 and hence be of the form... y=-1/5(x) + c substitutimng the values of the ordered pair given 9= -1/5(5) + c 9= -1 + c adding 1 to both sides 10 = c now you can complete the equation of the perpindicular y = 10 - 1/5(x)

OpenStudy (blacksteel):

Okay, in order for two lines to be perpendicular, the slopes of the lines must be negative reciprocals of one another. The line you are given is y = 8 - 5x, which has a slope of -5. Then your new line must have a slope of 1/5. Thus, y = 1/5x + k. Now, we determine k. Since the line passes through (5,9), we know that f(5) = 1 + k = 9. Then k = 8. Thus your line is y = 1/5x + 8. Also, the above answer is close but wrong; he didn't move the x term to the right side before differentiating.

OpenStudy (anonymous):

Black steel you are correct. A brutal error on my behalf!

OpenStudy (blacksteel):

Haha, we all make them!